Oxygen consumed by a living organism in aerobic re
Answer #1

The balanced reaction is

C6H12O6+ NH3 +1.5O2---àC5H9NO4 +CO2 +3H2O

Moles of C6H12O6= 100 mol/day,

NH3= 120 mol/day

And O2= 110 moles/day

Molar ratio of C6H12o6: NH3 : O2= 1:1:1.5

Actual ratio = 100:120:110 = 1:1.2:1.1

Limiting reactants is Oxygen.

Since all the limiting reactant is consumed, given

Rq= moles of CO2 produced/ mole of O2 consumed = 0.45

Moles of CO2 produced = 0.45*110=49.5 moles

Moles of excess reactant consumed = 49.5 moles

Products contains :CO2= 49.5 moles/day, C6H9NO4= 49.5 moles/day, H2O= 3*49.5= 148.5 moles/day

NH3= 120-49.5= 70.5 moles/day, O2=110-49.5*1.5=35.75moles/day

Fractional conversion of NH3 =100*49.5/120 = 41.25%

Fractional conversion of C6H12O6= 100*49.5/100=49.5%